OpenStudy (anonymous):

CAN SOMEONE PPPLLLZZZ HELP ME!!!!!! what is the value of √9y^4+12√y^6-3√y^6-√y^4 A. 8y^2+9y^3 B. 11y^2 C. 11y^5 D. 2y^2+9y^3

OpenStudy (anonymous):

\[\sqrt{9y ^{4}}=3y ^{2}\] \[12\sqrt{y ^{6}}=12y ^{3}\] \[-3\sqrt{y ^{6}}=-3y ^{3}\] \[-\sqrt{y ^{4}}=-y ^{2}\] add them

OpenStudy (anonymous):

so it would be 11y^5

OpenStudy (anonymous):

no!! 2y^2+9y^3 :)

OpenStudy (anonymous):

but that doesnt make sense if u added -√4^4 and -y^2 how do u get that????

OpenStudy (anonymous):

i don't understand what you asked

OpenStudy (anonymous):

how did you get that answer

OpenStudy (anonymous):

3y^2-y^2=2y^2 and 12y^3-3y^3=9y^3 so answer is 2y^2+9y^3